The sum of radii of two circles is 140 cm and the difference of their circumference is 88 cm find the ratio between their areas

Grade:10

1 Answers

Sandeep Bhamoo
5 years ago
R1+R2 = 140, 2π(R1-R2)=88. R1-R2=14 on dolving both equation R1=77, R2=63. Ratio of area= R1²/R2²= (77/63)²= 121/81. Pleae do approve my answer if correct